# Point modules over the universal enveloping algebras of color Lie algebras and overlapping normal elements

### 01 Apr 2026 v1

@@ 4.1. Overlapping normal elements and truncated point modules
Let $A$ be a $\mathbb{Z}$-graded $k$-algebra (not necessarily connected, $\mathbb{N}$-graded, finitely presented, or generated in degree $1$) and g=xyuyxg=xy-uyx an overlapping normal element of $A$ with degree $n$. Then, for an integer $r\geq 1$, the following equations hold;
@@
See [10, Lemma 2.3]. ∎
The display of an overlapping normal element $g$ of $A$ is not necessarily unique, that is, we may have displays $g=xy-uyx=x^{\prime}y^{\prime}-u^{\prime}y^{\prime}x^{\prime}$ where $x,x^{\prime}\in A_{1}$, $y,y^{\prime}\in A_{n-1}$, and $u,u^{\prime}\in k^{\times}$. However, when $g$ is displayed by g=xyuyxg=xy-uyx, a $g$-torsionfree truncated point module over $A$ has a nice relationship to $x$.
@@
Let $A$ be a connected $\mathbb{N}$-graded finitely presented $k$-algebra generated in degree $1$, g=xyuyxg=xy-uyx an overlapping normal element of $A$ with degree $n$ of $A$, and $P=\bigoplus_{i=0}^{d}km_{i}$ a truncated $g$-torsionfree point module of length $d+1$ over $A$ where $d\geq 2n-1$. Then $xm_{i}\neq 0$ holds for $0\leq i\leq d-1$.
@@
Proposition 4.2 shows the nonexistence of a long truncated $g$-torsionfree point module over $A$ where $g$ is an overlapping normal element of $A$.
@@
Let $A$ be a connected $\mathbb{N}$-graded finitely presented $k$-algebra generated in degree $1$ and g=xyuyxg=xy-uyx an overlapping normal element of $A$ with degree $n$. For $d\geq 2n-1$, $A$ has no a truncated $g$-torsionfree point module of length $d+1$ over $A$.
@@
The notation in the proof of Proposition 4.2 is retained. Let $\{r_{i}\}_{i=1}^{n+1}$ be a sequence defined by $r_{i}=\dfrac{q_{i+n-2}}{p_{i+n-2}p_{i+n-3}\cdots p_{i}}$ and $P=\bigoplus_{i=0}^{d}km_{i}$ a truncated $g$-torsionfree point module of length $d+1$ over $A$ where $d\geq 2n-1$. The definition of an overlapping normal element of $A$ gives the following equations;
@@
Remark that, since a truncated $g$-torsionfree point module over $A$ is not a $g$-torsionfree module over $A$, Theorem 4.3 is not a corollary of Theorem 3.4.
@@ 4.2. Point modules over the universal enveloping algebras of color Lie algebras
Hence, $g$ is an overlapping normal element of $U(L)$ with $\deg g=n_{L}\geq 2$. Therefore, for $d\geq 2n_{L}-1$, the set of isomorphism classes of truncated point modules of length $d+1$ over $U(L)$ corresponds one-to-one to the one of $U(L)/\langle g\rangle$ by Proposition 2.12 and Theorem 4.3. By the similar proof of [9, 2.2.14], we have $U(L)/\langle g\rangle\cong U(L/\langle g\rangle)$. Also, clearly $d\geq 2n_{L}-1\geq 2n_{L/\langle g\rangle}-1$ and $S_{\varepsilon}(L_{\mathbf{1}})\cong S_{\varepsilon}((L/\langle g\rangle)_{\mathbf{1}})$. By induction hypothesis, we have conclusion. The second claim is follows from Theorem 4.4. ∎
Note that the second claim of Theorem 4.5 can also be followed by Theorem 3.4. On the other hands, the following corollary is only followed by Theorem 4.3.
@@
For a scheme $X$, the set of closed points of $X$ is denoted by $\mathrm{Cl}(X)$. By Theorem 4.5, we obtain an equation $\mathrm{Cl}(\Gamma_{2n_{L}-1}(U(L)))=\mathrm{Cl}(\Gamma_{2n_{L}-1}(S_{\varepsilon}(L_{\mathbf{1}})))$. Also, for a normal element $g$ of $U(L)$ where $g\in h(L^{n_{L}}_{\mathbf{1}})$, Propositions 2.12 and 2.13 provide $\mathrm{Cl}(\Gamma_{2n_{L}-2}(U(L/\langle g\rangle)))\subset\mathrm{Cl}(\Gamma_{2n_{L}-2}(U(L)))$. Inductively, we obtain
@@
Then, by [22, 11.2.Q.Exercise] and [5] (also, see [27]), $\mathrm{pr}_{1,2n_{L}-2}^{(2n_{L}-1)}:\Gamma_{2n_{L}-1}(U(L))\rightarrow\Gamma_{2n_{L}-2}(U(L))$ is injective on points. By [2, Proposition 3.6(ii)], $\mathrm{pr}_{1,2n_{L}-2}^{(2n_{L}-1)}$ is a closed immersion and identify a closed subscheme $E\subset\Gamma_{2n_{L}-2}(U(L))$. Also, $\sigma:E\rightarrow\Gamma_{2n_{L}-2}(U(L))$ is defined as $(p_{1},\cdots,p_{2n_{L}-2})\mapsto(p_{2},\cdots,p_{2n_{L}-1})$ where $(p_{1},\cdots,p_{2n_{L}-1})\in\Gamma_{2n_{L}-1}(U(L))$ is the point laying over $(p_{1},\cdots,p_{2n_{L}-2})$ by [2, Proposition 3.7(i)]. By similar discussion in above, $\sigma(E)\subset E$ is obtained.
@@
It follows that $\partial_{2}(w)=\partial_{2}\pi_{2}f_{2}(w)=\partial_{2}\pi_{2}(z)$ and then $\partial_{2}(P_{2})\subset\partial_{2}(Q^{\prime})$ is obtained. This contradicts the minimality of $P_{\bullet}$. Hence, $P_{2}$ is isomorphic to a direct summand of $\operatorname*{\mathchoice{\Xdbigoplus}{\Xdbigoplus}{\Xtbigoplus}{\Xtbigoplus}}\limits_{|b|_{\mathbb{Z}}+|b^{\prime}|_{\mathbb{Z}}<2n_{L},\,(b,b^{\prime})\in B_{2}}U(L)(-|b|_{\mathbb{Z}}-|b^{\prime}|_{\mathbb{Z}})$. This means that $I$ is generated by relations with degree $\leq 2n_{L}-1$ or less by [20, Lemma 2.1.3]. Therefore, by [2, Proposition 3.7(ii)], for $r\geq 2n_{L}$, $\mathrm{pr}_{1,2n_{L}-1}^{(r)}:\Gamma_{r}(U(L))\rightarrow E$ is isomorphism and then $\Gamma_{U(L)}=\displaystyle{\lim_{\leftarrow}}\Gamma_{j}(U(L))\cong E$. ∎

### 11 Apr 2026 v2

@@ 4.1. $q^{\prime}$-Heisenberg normal elements and truncated point modules
Let $A$ be a $\mathbb{Z}$-graded $k$-algebra (not necessarily connected, $\mathbb{N}$-graded, finitely presented, or generated in degree $1$) and g=xyuyxg=xy-uyx a $q^{\prime}$-Heisenberg normal element of $A$ with degree $n$. Then, for an integer $r\geq 1$, the following equations hold;
@@
See [12, Lemma 2.3]. ∎
The display of a $q^{\prime}$-Heisenberg normal element $g$ of $A$ is not necessarily unique, that is, we may have displays $g=xy-uyx=x^{\prime}y^{\prime}-u^{\prime}y^{\prime}x^{\prime}$ where $x,x^{\prime}\in A_{1}$, $y,y^{\prime}\in A_{n-1}$, and $u,u^{\prime}\in k^{\times}$. However, when $g$ is displayed by g=xyuyxg=xy-uyx, a $g$-torsionfree truncated point module over $A$ has a nice relationship to $x$.
@@
Let $A$ be a connected $\mathbb{N}$-graded finitely presented $k$-algebra generated in degree $1$, g=xyuyxg=xy-uyx a $q^{\prime}$-Heisenberg normal element of $A$ with degree $n$ of $A$, and $P=\bigoplus_{i=0}^{d}km_{i}$ a truncated $g$-torsionfree point module of length $d+1$ over $A$ where $d\geq 2n-1$. Then $xm_{i}\neq 0$ holds for $0\leq i\leq d-1$.
@@
Proposition 4.2 shows the nonexistence of a long truncated $g$-torsionfree point module over $A$ where $g$ is a $q^{\prime}$-Heisenberg normal element of $A$.
@@
Let $A$ be a connected $\mathbb{N}$-graded finitely presented $k$-algebra generated in degree $1$ and g=xyuyxg=xy-uyx a $q^{\prime}$-Heisenberg normal element of $A$ with degree $n$. For $d\geq 2n-1$, $A$ has no a truncated $g$-torsionfree point module of length $d+1$ over $A$.
@@
The notation in the proof of Proposition 4.2 is retained. Let $\{r_{i}\}_{i=1}^{n+1}$ be a sequence defined by $r_{i}=\dfrac{q_{i+n-2}}{p_{i+n-2}p_{i+n-3}\cdots p_{i}}$ and $P=\bigoplus_{i=0}^{d}km_{i}$ a truncated $g$-torsionfree point module of length $d+1$ over $A$ where $d\geq 2n-1$. The definition of a $q^{\prime}$-Heisenberg normal element of $A$ gives the following equations;
@@
Remark that, since a truncated $g$-torsionfree point module over $A$ is not a $g$-torsionfree module over $A$, Theorem 4.3 is not a corollary of Proposition 3.7.
@@ 4.2. Point modules over the universal enveloping algebras of color Lie algebras
Hence, $g$ is a $q^{\prime}$-Heisenberg normal element of $U(L)$ with $\deg g=n_{L}\geq 2$. Therefore, for $d\geq 2n_{L}-1$, the set of isomorphism classes of truncated point modules of length $d+1$ over $U(L)$ corresponds one-to-one to the one of $U(L)/\langle g\rangle$ by Proposition 2.13 and Theorem 4.3. By the similar proof of [11, 2.2.14], we have $U(L)/\langle g\rangle\cong U(L/\langle g\rangle)$. Also, clearly $d\geq 2n_{L}-1\geq 2n_{L/\langle g\rangle}-1$ and $S_{\varepsilon}(L_{\mathbf{1}})\cong S_{\varepsilon}((L/\langle g\rangle)_{\mathbf{1}})$. By induction hypothesis, we have conclusion. The second claim is follows from Theorem 4.4. ∎
Note that the second claim of Theorem 4.5 can also be followed by Proposition 3.7. On the other hands, the following corollary is only followed by Theorem 4.3.
@@
For a scheme $X$, the set of closed points of $X$ is denoted by $\mathrm{Cl}(X)$. By Theorem 4.5, we obtain an equation $\mathrm{Cl}(\Gamma_{2n_{L}-1}(U(L)))=\mathrm{Cl}(\Gamma_{2n_{L}-1}(S_{\varepsilon}(L_{\mathbf{1}})))$. Also, for a normal element $g$ of $U(L)$ where $g\in h(L^{n_{L}}_{\mathbf{1}})$, Propositions 2.13 and 2.14 provide $\mathrm{Cl}(\Gamma_{2n_{L}-2}(U(L/\langle g\rangle)))\subset\mathrm{Cl}(\Gamma_{2n_{L}-2}(U(L)))$. Inductively, we obtain
@@
Then, by [25, 11.2.Q.Exercise] and [6] (also, see [30]), $\mathrm{pr}_{1,2n_{L}-2}^{(2n_{L}-1)}:\Gamma_{2n_{L}-1}(U(L))\rightarrow\Gamma_{2n_{L}-2}(U(L))$ is injective on points. By [2, Proposition 3.6(ii)], $\mathrm{pr}_{1,2n_{L}-2}^{(2n_{L}-1)}$ is a closed immersion and identify a closed subscheme $E\subset\Gamma_{2n_{L}-2}(U(L))$. Also, $\sigma:E\rightarrow\Gamma_{2n_{L}-2}(U(L))$ is defined as $(p_{1},\cdots,p_{2n_{L}-2})\mapsto(p_{2},\cdots,p_{2n_{L}-1})$ where $(p_{1},\cdots,p_{2n_{L}-1})\in\Gamma_{2n_{L}-1}(U(L))$ is the point laying over $(p_{1},\cdots,p_{2n_{L}-2})$ by [2, Proposition 3.7(i)]. By similar discussion in above, $\sigma(E)\subset E$ is obtained.
@@
It follows that $\partial_{2}(w)=\partial_{2}\pi_{2}f_{2}(w)=\partial_{2}\pi_{2}(z)$ and then $\partial_{2}(P_{2})\subset\partial_{2}(Q^{\prime})$ is obtained. This contradicts the minimality of $P_{\bullet}$. Hence, $P_{2}$ is isomorphic to a direct summand of $\operatorname*{\mathchoice{\Xdbigoplus}{\Xdbigoplus}{\Xtbigoplus}{\Xtbigoplus}}\limits_{|b|_{\mathbb{Z}}+|b^{\prime}|_{\mathbb{Z}}<2n_{L},\,(b,b^{\prime})\in B_{2}}U(L)(-|b|_{\mathbb{Z}}-|b^{\prime}|_{\mathbb{Z}})$. This means that $I$ is generated by relations with degree $\leq 2n_{L}-1$ or less by [23, Lemma 2.1.3]. Therefore, by [2, Proposition 3.7(ii)], for $r\geq 2n_{L}$, $\mathrm{pr}_{1,2n_{L}-1}^{(r)}:\Gamma_{r}(U(L))\rightarrow E$ is isomorphism and then $\Gamma_{U(L)}=\displaystyle{\lim_{\leftarrow}}\Gamma_{j}(U(L))\cong E$. ∎